Browns, a local bakery, is worried about increased costs particularly energy. Last years records can provide a fairly good   Last Year   Now
Production (dozen)   1,500   1,500
Labor (hours)   350   325
Capital Investment ($)   16,000   18,500
Energy (BTU)   3,000   2,800

Please answer and show all work.
Answer #1

Energy productivity last year = (1500*12)/3000 =6 loaves/BTU

Energy productivity Now = (1500*12)/2800 =6.43 loaves/BTU

Percentage change =[ (6.43-6)/6]*100% =0.07167 *100%

ercentage change =7.17%

KINDLY RATE THE ANSWER AS THUMBS UP. THANKS A LOT
Answer #2
But why 1500 multiply with 12 as there is no explanation of calculating the month
source: Book
answered by: anonymous